# Latent Action Learning Requires Supervision in the Presence of Distractors

### v2

@@ 1 Introduction
Recently, a new wave of approaches based on latent action learning has emerged (Edwards et al., 2019), demonstrating superior pre-training efficiency on datasets without action labels in large-scale robotics (Ye et al., 2024; Chen et al., 2024; Cui et al., 2024; Bruce et al., 2024) and reinforcement learning (Schmidt & Jiang, 2023). Latent Action Models (LAM) infer latent actions between successive observations, effectively compressing observed changes. Under certain conditions, latent actions can even rediscover the ground truth action space (Schmidt & Jiang, 2023; Bruce et al., 2024). After training, LAM can be utilized for imitation learning on latent actions to obtain useful behavioral priors. For example, LAPA (Ye et al., 2024) showed that latent action learning can be used to pre-train large model on only human manipulation videos, and despite the huge cross-embodiment gap, still outperform OpenVLA (Kim et al., 2024b), which was pre-trained on expert in-domain data with available action labels.
@@
In this work we empirically investigate the effect of action-correlated distractors on latent action learning using Distracting Control Suite (Stone et al., 2021). We demonstrate that naive latent action learning based on quantization and reconstruction objectives, such as LAPO (Schmidt & Jiang, 2023), struggle in the precense of distractors (see Section 4). We propose LAOM, a simple LAPO modification that improve the quality of latent actions by 8x, as measured by linear probing, and double the downstream performance (see Figure 7). However, even after this, the resulting performance is only slightly better than simple Behavioral Cloning on available ground-truth actions. Thus, as our core contribution, we show that providing supervision with a small number, as little as $2.5$% of the complete dataset, of action labels during LAOM training improves the downstream performance by 4.3x on average (see Section 5), outperforming all baselines (see Figure 7). Our findings suggest that the pipeline used in most current work (Ye et al., 2024; Cui et al., 2024; Chen et al., 2024) to first learn LAM and only then decode to ground-truth actions is suboptimal when distractors are present, as with supervision better result can be achieved using the same budget of actions labels. In addition, we show that latent action learning with supervision generalizes better in contrast to approaches based on inverse dynamics models (Baker et al., 2022; Zhang et al., 2022a; Zheng et al., 2023) but does not learn control-endogenous minimal state (Lamb et al., 2022).
@@ 2 Preliminaries
Latent action learning. Latent action learning approaches (Edwards et al., 2019; Schmidt & Jiang, 2023; Chen et al., 2024; Cui et al., 2024; Ye et al., 2024) aim to infer latent actions $z_{t}$ such that they are maximally informative about each observed transition $(o_{t},o_{t+1})$ while being minimal. After the latent action model (LAM) is pre-trained, we can train policies to imitate latent actions on full data to obtain useful behavioral priors. Finally, small decoder heads can be learned from latent to real actions of domain of interest.
We base our work on LAPO (Schmidt & Jiang, 2023), which is used in recent work (Chen et al., 2024; Cui et al., 2024; Ye et al., 2024). LAPO uses two models in combination to infer latent actions. First is inverse dynamics model (IDM), which is given two consecutive observations predicts latent action $z_{t}\sim p_{\text{IDM}}(\cdot|o_{t},o_{t+1})$. Second is forward dynamics model (FDM), which observes current observation and latent action, and predicts the next observation $\hat{o}_{t+1}\sim p_{\text{FDM}}(\cdot|o_{t},z_{t})$. Both models are trained jointly to minimize the next observation prediction loss $\|\hat{o}_{t+1}-o_{t+1}\|^{2}$. We illustrate the model architecture in Figure 2.
Given the information bottleneck on latent actions, e.g. quantization via the VQ-VAE (Van Den Oord et al., 2017), IDM cannot simply copy the next observation into the FDM as is, so it will be forced to compress and encode the difference between observations to be most predictive of the next observation. Without the distractors, through simplicity bias (Shah et al., 2020), the latent actions will recover the ground truth actions as they are most predictive of the dynamics. However, due to the presence of distractors, it may be not true for real-world data. In this work, we empirically examine how well current LAM can recover true actions in such circumstances.
Figure 3: We visualize the environments from Distracting Control Suite used in our work. Top row: without any distractors, identical to the original DeepMind Control Suite. Bottom row: with distractors, which consists of dynamic background videos, agent color change and camera shaking.
@@ 3 Experimental Setup
As currently existing benchmarks with distractors (Stone et al., 2021; Ortiz et al., 2024) are not yet solved, we collect new datasets with custom difficulty, based on Distracting Control Suite (DCS) (Stone et al., 2021). DCS uses dynamic background videos, camera shaking and agent color change as distractors (see Figure 3 for visualization). The complexity is determined by the number of videos as well as the scale for the magnitude of the camera and the color change. We empirically found that using 60 videos and a scale of 0.1 is the hardest setting when BC can still recover expert performance. We collect datasets with five thousand trajectories for four tasks: cheetah-run, walker-run, hopper-hop and humanoid-walk, listed in the order of increasing difficulty. See Appendix C for additional details.
@@
Evaluation. To access the quality of the latent actions, we use two methods. First, we follow the approach of Zhang et al. (2022b) and use linear probing (Alain, 2016), which is a common technique used to evaluate the quality of learned representations by training a simple linear classifier or regressor on top the representations. Since we include ground truth actions in our datasets for debugging purposes, we train linear probes to predict them from latent actions simultaneously with the main method, e.g. LAPO (Schmidt & Jiang, 2023). We do not pass the gradient through the latent actions, so this does not affect the training. Second, following the most commonly used three-stage pipeline (Schmidt & Jiang, 2023; Chen et al., 2024; Ye et al., 2024), we first pre-train LAM, then train BC model to predict latent actions on the full dataset, and finally, we reveal a small number of labeled trajectories to train a small two-layer MLP decoder from latent to real actions. Using this decoder, we then evaluate the resulting agent in the environment for 25 episodes. To access scaling properties with different budgets of real actions, similar to Schmidt & Jiang (2023), we repeat this process for a variable number of labeled trajectories, from 2 to 128. All experiments are averaged over three random seeds.
@@
Baselines. We use BC on true actions as our main baseline, since the main goal of latent action learning is to pre-train useful behavioral policies (Edwards et al., 2019; Schmidt & Jiang, 2023), which can be achieved by recovering true actions as accurately as possible. We use it in two ways. First, we try to get the best performance for each full dataset with true actions to use the final return for normalization. With such normalization, we can quantify how much performance we have recovered compared to if we had access to a fully action-labeled dataset. Second, we train BC from scratch on the same number of labels available to LAM, to evaluate the benefit of pre-training on large unlabeled data. Our last baseline is IDM, as it remains one of the most successful and simplest approaches to learn from action-free data at scale (Baker et al., 2022; Zhang et al., 2022a; Zheng et al., 2023). For additional details, see Appendix E.
@@
On hyperparameters tuning. We tune the hyperparameters based on online performance for BC, on MSE to real actions on the full dataset for IDM, and on final linear probe MSE to real actions for latent action learning. In more practical tasks, we usually do not have this luxury, but since we are interested in estimating the upper bound performance of each method in a controlled setting, we believe that it is appropriate. For exact hyperparameters see Appendix G.
@@ 4 Latent Action Learning Struggle in the Presence of Distractors
Quantization hinders latent action learning. To validate our implementation, we first measured performance on distractor-free datasets, which should not cause any difficulty. Contrary to previous research (Schmidt & Jiang, 2023; Chen et al., 2024; Ye et al., 2024; Bruce et al., 2024), we found that commonly used latent action quantization during training significantly hindered the resulting latent action quality. We initially hypothesized that the problem might be with the VQ-VAE used for quantizing. In conversation with Schmidt & Jiang (2023) we confirmed that VQ-VAE is indeed susceptible to codebook collapse and requires extensive tuning. We tried the more modern FSQ (Mentzer et al., 2023), which has already been used successfully in RL (Scannell et al., 2024) and does not suffer from codebook collapse. Unfortunately, even after tuning, we were unable to improve the results significantly, so we simply removed it. To our surprise, this resulted in a large positive improvement (see Figure 4), but only for datasets without distractors, while with distractors the action quality remained at almost the same level.
@@
5 Latent Action Learning Requires Supervision in the Presence of Distractors
@@
Supervision significantly increases downstream performance. Despite the fact that existing approaches (Schmidt & Jiang, 2023; Ye et al., 2024; Chen et al., 2024) pre-train LAM without true actions, in practice we still need to have some number of labels to learn the action decoder as last stage. We reuse these labels to provide supervision by linearly predicting them from latent actions during LAOM training (see Figure 2 for the final architecture). We plot the resulting downstream performance for each environment in Figure 7 and summarize in Figure 1. As can be seen, LAOM+supervision outperforms all baselines and scales better with a larger budget of real actions. It achieves an average normalized score of 0.44, i.e. it recovers almost half the performance of BC with access to the full dataset of true actions, while using only 2.5% of the action labels. Importantly, all methods have access to exactly the same number of action labels, differing only in how they use them. We provide results for distractor-free data in Appendix F.
@@ 6 Related Work
In contrast to previous work (Schmeckpeper et al., 2020; Baker et al., 2022; Zheng et al., 2023), we show that while IDM is a strong baseline in setups without distractors (see Figure 12 in Appendix F), it generalizes poorly when distractors are present. Our results show that when a small number of action labels are available, it is much better to combine IDM and latent action learning to achieve much stronger performance and generalization (see Figure 7), suggesting that for web-scale data (Baker et al., 2022; Zhang et al., 2022a) our approach may be better than simple IDM re-labeling.
Latent action learning. To our knowledge, Edwards et al. (2019) was the first to propose the task of recovering latent actions and imitating latent policies from observation, with limited success on simple problems. However, the original objective had scalability issues (Struckmeier & Kyrki, 2023). LAPO (Schmidt & Jiang, 2023) greatly simplified the approach, removed scalability barriers, and for the first time achieved high success on the hard, procedurally generated ProcGen benchmark (Cobbe et al., 2020). Latent action learning was further scaled by Bruce et al. (2024); Cui et al. (2024); Ye et al. (2024); Chen et al. (2024) to larger models, data, and harder, more diverse robotics domains.
In contrast to our work, all the mentioned approaches (Schmidt & Jiang, 2023; Ye et al., 2024; Cui et al., 2024; Chen et al., 2024) use data without distractors, where all changes in dynamics are mainly explained by ground truth actions only. As we show in our work (see Section 4), naive latent action learning does not work in the presence of distractors. Although we propose improvements that double the performance, it is not enough (see Figure 6). Providing supervision with a small number of action labels during LAM training significantly improves performance (see Figure 1), suggesting that the pipeline used in most current work (Ye et al., 2024; Cui et al., 2024; Chen et al., 2024) to first learn LAM and only then decode to ground-truth actions is suboptimal.
@@
7 Conclusion
In this work, we empirically investigated the effect of action-correlated distractors on latent action learning. We showed that LAPO struggles to learn latent actions useful for pre-training. Although we proposed LAOM, a simple modification of LAPO, which doubled performance, it did not fully close the gap with the distractor-free setting. Crucially, we found that even minimal supervision - reusing as little as 2.5% of the dataset’s ground-truth action labels during latent action learning significantly improved downstream performance, challenging the conventional pipeline of first pre-training LAM and only then decoding from latent to real actions. Our findings suggest that integrating supervision is essential for robust latent action learning in real-world scenarios, paving the way for unlocking the vast amounts of video data available on the web for embodied AI. We discuss the limitations of our work in the Appendix A.
Impact Statement
Appendix A Limitations
@@
Appendix B Additional Related Work
@@
Appendix C Data Collection
@@
Appendix D Implementation Details
@@
For all experiments we use cosine learning late schedule with warmup. We will publicly release the code, all configs and all Weights&Biases logs after the review. For hyperparameters see Appendix G.
@@
Appendix E Evaluation Details
We outline the evaluation procedures used in our experiments for each method. First, we review the general setup. For each environment, we have a large dataset without action labels, with and without distractors. To decode the learned latent actions to ground truth for evaluation, we allow a small amount of action labeled data, in line with previous work (Schmidt & Jiang, 2023; Ye et al., 2024). We sample it once from the existing dataset, revealing true actions, to ensure that all methods are on equal conditions. We use identical backbones where possible, and try our best to make all methods equal in the number of trainable weights. For hyperparameters, see Appendix G. We report the scores achieved by BC trained on datasets with all actions revealed in Table 4. We use these for normalization in all our experiments.

### v3

@@ 1 Introduction
Figure 2: Visualization of the environments from the Distracting Control Suite (DCS) used in our work. Top row: without any distractors, identical to the original DeepMind Control Suite. Bottom row: with distractors, which consists of dynamic background videos, agent color change and camera shaking. See Section 3 for additional details.
Recently, a new wave of approaches based on latent action learning has emerged (Edwards et al., 2019), demonstrating superior pre-training efficiency on datasets without action labels in large-scale robotics (Ye et al., 2024; Chen et al., 2024b, a; Cui et al., 2024; Bruce et al., 2024) and reinforcement learning (Schmidt & Jiang, 2023). Latent Action Models (LAM) infer latent actions between successive observations, effectively compressing observed changes. Under certain conditions, latent actions can even rediscover the ground truth action space (Schmidt & Jiang, 2023; Bruce et al., 2024). After training, LAM can be utilized for imitation learning on latent actions to obtain useful behavioral priors. For example, LAPA (Ye et al., 2024) showed that latent action learning can be used to pre-train large model on only human manipulation videos, and despite the huge cross-embodiment gap, still outperform OpenVLA (Kim et al., 2024b), which was pre-trained on expert in-domain data with available action labels.
@@
In this work we empirically investigate the effect of action-correlated distractors on latent action learning using Distracting Control Suite (Stone et al., 2021). We demonstrate that naive latent action learning based on quantization and reconstruction objectives, such as LAPO (Schmidt & Jiang, 2023), struggle in the precense of distractors (see Section 4). We propose LAOM, a simple LAPO modification that improve the quality of latent actions by 8x, as measured by linear probing, and double the downstream performance (see Figure 8). However, even after this, the resulting performance is only slightly better than simple Behavioral Cloning on available ground-truth actions. Thus, as our core contribution, we show that providing supervision with a small number, as little as $2.5$% of the complete dataset, of action labels during LAOM training improves the downstream performance by 4.3x on average (see Section 5), outperforming all baselines (see Figure 8). Our findings suggest that the pipeline used in most current work (Ye et al., 2024; Cui et al., 2024; Chen et al., 2024b) to first learn LAM and only then decode to ground-truth actions is suboptimal when distractors are present, as with supervision better result can be achieved using the same budget of actions labels. In addition, we show that latent action learning with supervision generalizes better in contrast to approaches based on inverse dynamics models (Baker et al., 2022; Zhang et al., 2022a; Zheng et al., 2023) but does not learn control-endogenous minimal state (Lamb et al., 2022).
Figure 3: Overview of the latent action learning pipeline. In the first stage, the Latent Action Model (LAM) is pre-trained to infer latent actions between consecutive observations. In the second stage, the LAM is used to relabel the entire dataset with latent actions, which are then used for behavioral cloning. Finally, a decoder is trained to map from latent to true actions using a small number of labelled trajectories. In our work, we do not modify this pipeline in any way; we only examine the LAM architecture itself (see Figure 4).
@@ 2 Preliminaries
Latent action learning. Latent action learning approaches (Edwards et al., 2019; Schmidt & Jiang, 2023; Chen et al., 2024b; Cui et al., 2024; Ye et al., 2024) aim to infer latent actions $z_{t}$ such that they are maximally informative about each observed transition $(o_{t},o_{t+1})$ while being minimal. After the latent action model (LAM) is pre-trained, we can train policies to imitate latent actions on full data to obtain useful behavioral priors. Finally, small decoder heads can be learned from latent to real actions of domain of interest.
We base our work on LAPO (Schmidt & Jiang, 2023), which is used in recent work (Chen et al., 2024b; Cui et al., 2024; Ye et al., 2024). LAPO uses two models in combination to infer latent actions. First is inverse dynamics model (IDM), which is given two consecutive observations predicts latent action $z_{t}\sim p_{\text{IDM}}(\cdot|o_{t},o_{t+1})$. Second is forward dynamics model (FDM), which observes current observation and latent action, and predicts the next observation $\hat{o}_{t+1}\sim p_{\text{FDM}}(\cdot|o_{t},z_{t})$. Both models are trained jointly to minimize the next observation prediction loss $\|\hat{o}_{t+1}-o_{t+1}\|^{2}$. We illustrate the model architecture in Figure 4.
Given the information bottleneck on latent actions, e.g. quantization via the VQ-VAE (Van Den Oord et al., 2017), IDM cannot simply copy the next observation into the FDM as is, so it will be forced to compress and encode the difference between observations to be most predictive of the next observation. Without the distractors, through simplicity bias (Shah et al., 2020), the latent actions will recover the ground truth actions as they are most predictive of the dynamics. However, due to the presence of distractors, it may be not true for real-world data. In this work, we empirically examine how well current LAMs can recover true actions in such circumstances.
@@ 3 Experimental Setup
As currently existing benchmarks with distractors (Stone et al., 2021; Ortiz et al., 2024) are not yet solved, we collect new datasets with custom difficulty, based on Distracting Control Suite (DCS) (Stone et al., 2021). DCS uses dynamic background videos, camera shaking and agent color change as distractors (see Figure 2 for visualization). The complexity is determined by the number of videos as well as the scale for the magnitude of the camera and the color change. We empirically found that using 60 videos and a scale of 0.1 is the hardest setting when BC can still recover expert performance. We collect datasets with five thousand trajectories for four tasks: cheetah-run, walker-run, hopper-hop and humanoid-walk, listed in the order of increasing difficulty. See Appendix B for additional details.
@@
Evaluation. To access the quality of the latent actions, we use two methods. First, we follow the approach of Zhang et al. (2022b) and use linear probing (Alain, 2016), which is a common technique used to evaluate the quality of learned representations by training a simple linear classifier or regressor on top the representations. Since we include ground truth actions in our datasets for debugging purposes, we train linear probes to predict them from latent actions simultaneously with the main method, e.g. LAPO (Schmidt & Jiang, 2023). We do not pass the gradient through the latent actions, so this does not affect the training. Second, following the most commonly used three-stage pipeline (Schmidt & Jiang, 2023; Chen et al., 2024b; Ye et al., 2024), we first pre-train LAM, then train BC model to predict latent actions on the full dataset, and finally, we reveal a small number of labeled trajectories to train a small two-layer MLP decoder from latent to real actions (see Figure 3). Using this decoder, we then evaluate the resulting agent in the environment for 25 episodes. To access scaling properties with different budgets of real actions, similar to Schmidt & Jiang (2023), we repeat this process for a variable number of labeled trajectories, from 2 to 128. All experiments are averaged over three random seeds.
@@
Baselines. We use BC on true actions as our main baseline, since the main goal of latent action learning is to pre-train useful behavioral policies (Edwards et al., 2019; Schmidt & Jiang, 2023), which can be achieved by recovering true actions as accurately as possible. We use it in two ways. First, we try to get the best performance for each full dataset with true actions to use the final return for normalization. With such normalization, we can quantify how much performance we have recovered compared to if we had access to a fully action-labeled dataset. Second, we train BC from scratch on the same number of labels available to LAM, to evaluate the benefit of pre-training on large unlabeled data. Our last baseline is IDM, as it remains one of the most successful and simplest approaches to learn from action-free data at scale (Baker et al., 2022; Zhang et al., 2022a; Zheng et al., 2023). For additional details, see Appendix D.
@@
On hyperparameters tuning. We tune the hyperparameters based on online performance for BC, on MSE to real actions on the full dataset for IDM, and on final linear probe MSE to real actions for latent action learning. In more practical tasks, we usually do not have this luxury, but since we are interested in estimating the upper bound performance of each method in a controlled setting, we believe that it is appropriate. For exact hyperparameters see Appendix F.
@@ 4 Latent Action Learning Struggle in the Presence of Distractors
Quantization hinders latent action learning. To validate our implementation, we first measured performance on distractor-free datasets, which should not cause any difficulty. Contrary to previous research (Schmidt & Jiang, 2023; Chen et al., 2024b; Ye et al., 2024; Bruce et al., 2024), we found that commonly used latent action quantization during training significantly hindered the resulting latent action quality. We initially hypothesized that the problem might be with the VQ-VAE used for quantizing. In conversation with Schmidt & Jiang (2023) we confirmed that VQ-VAE is indeed susceptible to codebook collapse and requires extensive tuning. We tried the more modern FSQ (Mentzer et al., 2023), which has already been used successfully in RL (Scannell et al., 2024) and does not suffer from codebook collapse. Unfortunately, even after tuning, we were unable to improve the results significantly, so we simply removed it. To our surprise, this resulted in a large positive improvement (see Figure 5), but only for datasets without distractors, while with distractors the action quality remained at almost the same level.
@@
5 Latent Action Learning Requires Supervision
@@
Supervision significantly increases downstream performance. Despite the fact that existing approaches (Schmidt & Jiang, 2023; Ye et al., 2024; Chen et al., 2024b) pre-train LAM without true actions, in practice we still need to have some number of labels to learn the action decoder as last stage. We reuse these labels to provide supervision by linearly predicting them from latent actions during LAOM training (see Figure 4 for the final architecture). We plot the resulting downstream performance for each environment in Figure 8 and summarize in Figure 1. As can be seen, LAOM+supervision outperforms all baselines and scales better with a larger budget of real actions. It achieves an average normalized score of 0.44, i.e. it recovers almost half the performance of BC with access to the full dataset of true actions, while using only 2.5% of the action labels. Importantly, all methods have access to exactly the same number of action labels, differing only in how they use them. We provide results for distractor-free data in Appendix E.
@@ 6 Related Work
In contrast to previous work (Schmeckpeper et al., 2020; Baker et al., 2022; Zheng et al., 2023), we show that while IDM is a strong baseline in setups without distractors (see Figure 13 in Appendix E), it generalizes poorly when distractors are present. Our results show that when a small number of action labels are available, it is much better to combine IDM and latent action learning to achieve much stronger performance and generalization (see Figure 8), suggesting that for web-scale data (Baker et al., 2022; Zhang et al., 2022a) our approach may be better than simple IDM re-labeling.
Latent action learning. To our knowledge, Edwards et al. (2019) was the first to propose the task of recovering latent actions and imitating latent policies from observation, with limited success on simple problems. However, the original objective had scalability issues (Struckmeier & Kyrki, 2023). LAPO (Schmidt & Jiang, 2023) greatly simplified the approach, removed scalability barriers, and for the first time achieved high success on the hard, procedurally generated ProcGen benchmark (Cobbe et al., 2020). Latent action learning was further scaled by Bruce et al. (2024); Cui et al. (2024); Ye et al. (2024); Chen et al. (2024b, a) to larger models, data, and harder, more diverse robotics domains.
In contrast to our work, all the mentioned approaches (Schmidt & Jiang, 2023; Ye et al., 2024; Cui et al., 2024; Chen et al., 2024b, a) use data without distractors, where all changes in dynamics are mainly explained by ground truth actions only. As we show in our work (see Section 4), naive latent action learning does not work in the presence of distractors. Although we propose improvements that double the performance, it is not enough (see Figure 7). Providing supervision with a small number of action labels during LAM training significantly improves performance (see Figure 1), suggesting that the pipeline used in most current work (Ye et al., 2024; Cui et al., 2024; Chen et al., 2024b, a) to first learn LAM and only then decode to ground-truth actions is suboptimal.
@@
7 Limitations
@@
8 Conclusion
In this work, we empirically investigated the effect of action-correlated distractors on latent action learning. We showed that LAPO struggles to learn latent actions useful for pre-training. Although we proposed LAOM, a simple modification of LAPO, which doubled performance, it did not fully close the gap with the distractor-free setting. Crucially, we found that even minimal supervision - reusing as little as 2.5% of the dataset’s ground-truth action labels during latent action learning significantly improved downstream performance, challenging the conventional pipeline of first pre-training LAM and only then decoding from latent to real actions. Our findings suggest that integrating supervision is essential for robust latent action learning in real-world scenarios, paving the way for unlocking the vast amounts of video data available on the web for embodied AI. We discuss the limitations of our work in the Section 7.
Impact Statement
Appendix A Additional Related Work
@@
Appendix B Data Collection
@@
Appendix C Implementation Details
@@
For all experiments we use the cosine learning late schedule with warmup. For hyperparameters see Appendix F. We open-source the code at https://github.com/dunnolab/laom.
@@
Appendix D Evaluation Details
We outline the evaluation procedures used in our experiments for each method. First, we review the general setup. For each environment, we have a large dataset without action labels, with and without distractors. To decode the learned latent actions to ground truth for evaluation, we allow a small amount of action labeled data, in line with previous work (Schmidt & Jiang, 2023; Ye et al., 2024). We sample it once from the existing dataset, revealing true actions, to ensure that all methods are on equal conditions. We use identical backbones where possible, and try our best to make all methods equal in the number of trainable weights. For hyperparameters, see Appendix F. We report the scores achieved by BC trained on datasets with all actions revealed in Table 4. We use these for normalization in all our experiments.